how to find potential on a perticular point in d complex electric circuit??


perticularly whn resistors and capacitors both r connected

Grade:Upto college level

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Potential difference between two points in circuit is the energy lost by the charge in being transferred from one point to another. For example, potential difference between A and B is found with following formula;

VAB=VB-VA=∑ε-∑i.R

 

This formula shows the energy lost by charge moving from point A to point B.

If the direction of the current and current passing through the resistor are the same then we take i.R as "+", if they are in opposite directions then we take i.R as "-".

VAB=VB-VA= -ε2-(+i.R1+i.R2)

VCB=VB-VC= -ε3-(-i.R3)

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi